# Latent Beam Diffusion Models for Decoding Image Sequences

### v2

@@ Abstract
While diffusion models excel at generating high-quality images from text prompts, they struggle with visual consistency in image sequences. Existing methods generate each image independently, leading to disjointed narratives — a challenge further exacerbated in non-linear storytelling, where scenes must connect beyond adjacent frames. We introduce a novel beam search strategy for latent space exploration, enabling conditional generation of full image sequences with beam search decoding. Unlike prior approaches that use fixed latent priors, our method dynamically searches for an optimal sequence of latent representations, ensuring coherent visual transitions. As the latent denoising space is explored, the beam search graph is pruned with a cross-attention mechanism that efficiently scores search paths, prioritizing alignment with both textual prompts and visual context. Human and automatic evaluations confirm that BeamDiffusion outperforms other baseline methods, producing full sequences with superior coherence, visual continuity, and textual alignment. All data and code implementations will be released after publication.
@@ 1 Introduction
Image diffusion models [43, 9, 44] have made significant advancements in generating high-quality images. These models, especially when combined with text prompts, have shown great potential in producing detailed and contextually accurate visuals [35, 29, 38]. Despite the impressive results at generating individual images based on specific text prompts, diffusion models face challenges when it comes to creating coherent sequences of images. In most cases, each image is generated independently based on its own prompt, which does not naturally ensure visual continuity across a sequence of correlated prompts. This problem is further compounded when the narrative is non-linear, where a scene may be connected not only to the immediate prior scene but also to earlier scenes in the sequence. This dynamic can result in sequences that lack coherence and visual consistency over time [2].
@@
We propose to utilize beam search as a potential solution to this problem. Beam search is a well-established technique in text generation tasks like machine translation [14, 51, 17, 37], allowing the model to explore multiple possible paths during the generation process, evaluating different trajectories and refining its predictions. However, despite its success in other domains, it has been barely explored in this context, leaving its potential untapped. In full image sequence generation, beam search could help navigate the complex latent space, refining coherence by adjusting to multiple text prompts or adding semantic cues.
@@
In this work, we explore beam search for latent prior selection in image sequence generation. While previous research has used latent representations as priors to improve sequence consistency [2, 33], it remains unclear which specific representations yield the best overall sequence. We build on these ideas by introducing beam search to explore the latent space, enabling iterative refinement of image sequences. As illustrated in Figure 1(a), our approach leverages search in the latent space to identify the best sequence of images, minimizing the risk of falling into suboptimal or incoherent paths, while maintaining visual continuity across images. As the beam diffusion algorithm traverses the latent space, the number of hypothesis grow and the least probably search paths are pruned to keep a fixed number of candidate paths. To this end, we incorporate a cross-attention mechanism [19] that evaluates and scores candidate image paths based on their alignment with both the textual description of the next step and the visual context of previous steps. This provides a reliable quality measure for each path, allowing us to retain the best ones while pruning the least suitable, ensuring both adherence to instructions and visual continuity throughout the sequence.
@@
We evaluated the effectiveness of BeamDiffusion in three domains: the Recipes domain and in the out-of-domain DIY and Visual Storytelling. Results obtained with human annotations, automatic evaluations and LLM-as-a-judge, all showed that the proposed approach outperforms baseline methods for generated sequence quality, showcasing its adaptability to diverse contexts.
@@ 3 Generating Non-Linear Visual Sequences
Generating a full sequence of visually coherent images from textual scene descriptions is a challenging research problem. Given a sequence of steps $S=\{s_{1},s_{2},\dots,s_{L}\}$, where each sequence step $s_{j}$ is a textual description, our goal is to generate a corresponding sequence of images $\{I_{1},I_{2},\dots,I_{L}\}$ that maintain consistency across all steps of the sequence. While Latent Diffusion Models (LDMs) [35] can generate high-quality images in a compressed latent space, this latent space is designed to generate images independently, starting from a randomly initialized latent $z_{T}$.
LDMs encode an input image $I$ into thelatent space representation $z$ using an encoder $\mathcal{E}$ and reconstruct it through a decoder $\mathcal{D}$ [35]. The reverse diffusion process is modeled by a U-Net backbone [36], $\epsilon_{\theta}$, which incorporates a cross-attention mechanism [47] to condition the denoising iterations on an external input, e.g. the output embeddings of a text encoder $\tau_{\theta}(s)$. LDMs are trained to iteratively estimate a denoised variant of $z$, with the loss
$L_{\text{LDM}}=\mathbb{E}_{\mathcal{E}(I),s,\epsilon\sim\mathcal{N}(0,1),t}\left[\|\epsilon-\epsilon_{\theta}(z_{t},t,\tau_{\theta}(s))\|_{2}^{2}\right]$ (1)
This loss evidences the lack of awareness of other denoising processes, making it challenging to ensure consistency acrosscorrelated prompts or previous denoising processes. This is particularly challenging in non-linear narratives [42], where maintaining character identity, lighting, and environmental details are crucial. To address these challenges, prior work [32, 21, 45, 2, 33] reuse latent denoising iterations from earlier steps in the sequence, paving the way to a key question: how to determine the optimal set of denoising latents that maximizes coherence across different generations? The answer lies in a combinatorial search space, that BeamDiffusion can traverse with a search strategy to determine an optimal solution, as we will discuss next.
@@ 3.1 Beam Diffusion Model
To address the limitations of standard LDMs in generating coherent image sequences, we introduce the BeamDiffusion model. The proposed approach aims to enhance both the consistency and quality of image sequences by adopting a strategy inspired by beam search, but with a twist, it operates on latent denoising space. BeamDiffusion incorporates latent representations from previously generated images. These latent representations guide the generation of new candidates, ensuring that each new image aligns with the context of the evolving sequence. This approach allows visual elements from previous images to be carried over or referenced, maintaining contextual consistency and continuity, while also adapting to the next scene description.
In the remainder of this article, the term step will always refer to an element in the input sequence; and the term iteration will refer to one denoising iteration of the diffusion model.
3.1.1 Latent Space Exploration with Diffusion Beams
@@
Algorithm 1 BeamDiffusion Algorithm
1:  Input parameters: $\{s_{1},\ldots,s_{N}\}$: sequence of steps; $r$: number of random seeds; $m$: steps back; $w$: beam width.
2:  $\hat{B}_{1}\leftarrow\{\text{SD}(s_{1},rnd)\}^{r}$triangleright Beam initialization with $r$ different images.
3:  for $j=2$ to $L$ do
4:     for all beam $b\in\hat{B}_{j-1}$ do
5:        $\mathcal{Z}_{j}\leftarrow$$\{\hat{B}_{j-n},\ldots,\hat{B}_{j-1}\}$triangleright Read denoising latents from previous steps.
6:        for all latent $z\in\mathcal{Z}_{j}$ do
7:           $I_{j}\leftarrow\text{SD}(s_{j},z)$triangleright Generate candidate a image.
8:           $\hat{B}_{j-1}\leftarrow\hat{B}_{j-1}\bigcup{I_{j}}$triangleright Extend beam with candidate image.
9:        end for
10:     end for
11:     $\text{score}(I_{j})\leftarrow\varphi(I_{j}\in\hat{B}_{j})$triangleright Score the new image of the beam.
12:     $B_{j}\leftarrow\arg\max_{|B_{j}|=w}\text{score}(\hat{B}_{j})$triangleright Prune beams and keep top $w$ beams.
13:  end for
14:  return  best beam $B^{*}\leftarrow\arg\max_{B_{L}}\text{score}(B_{L})$
@@
Inspired by beam search [50], which maintains and extends the most promising candidates at each sequence step, we introduce diffusion beams to structure sequential image generation. At each step $j$, the set of candidate images $\mathcal{I}_{j}$ is used to compute the set of candidate diffusion beams $\hat{B}_{j}$: for each step $j$ we consider all the beams existing in step $j-1$ and extend each one of such $b$ beams $\hat{B}_{j-1}^{(b)}$ by pairing it with a new candidate image $I_{j}^{(k)}\in\mathcal{I}_{j}$. The resulting beams are defined as
@@
3.1.2 Pruning Diffusion Beam Paths
@@
The contrastive classifier $\varphi(I_{j}|\{s_{1},\ldots,s_{j}\},\{I_{1},\ldots,I_{j-1}\})$ is trained as a next image prediction task to estimate how well align an image is with the previously sequence of steps and images. For a given sequence $t$ and sequence step $j$, the ground-truth label $l_{t,j}$ indicates whether image $I_{j}$ belongs to the sequence given its prompt $s_{j}$ and context $B_{j-1}^{(b)}$. Negative examples are hallucinated from other sequences. The model minimizes the cross-entropy loss encouraging the model to promote candidates that best match the sequence’s semantic and visual dynamics. See appendix A for implementation details.
@@ 4 Experimental Setup
In this section, we describe the experimental setup to evaluate BeamDiffusion’s performance in generating coherent image sequences. We outline the datasets used, identify other baseline methods, and explain the human, automatic and LLM-as-a-judge (Gemini [34]) evaluations focused on semantic and visual consistency across generated sequences.
@@
We used a dataset consisting of publicly available manual tasks in the Recipes and DIY domains [2], as well as the Visual Storytelling (VIST) dataset [13]. Recipes serve as the in-domain dataset, while DIY tasks and VIST are out-of-domain. These out-of-domain datasets allow us to assess the generalizability of our model across different task domains. More details can be found in Appendix B.2.
@@
To evaluate the performance of our image sequence generation method, we compared BeamDiffusion with SD2.1 [35] against two other competing approaches to generate image sequences: GILL [18] and StackedDiffusion [26]. Additionally, we considered two decoding strategies: Greedy and Nucleus Sampling [11]. As these decoding strategies are novel in the context of image generation, we provide a brief explanation of each. All methods used 50 denoising iterations.
In the Greedy approach, a new image is generated at each step using the first five latents from the previous image, with the image that scores highest overall being selected. Although this is similar to CoSeDlen=1 [33], it differs in that it relies exclusively on CLIP for selection without any learned module, whereas CoSeD incorporates a trained selection mechanism.
In contrast, Nucleus Sampling introduces diversity by sampling from the subset of images whose cumulative CLIP-assigned probability exceeds a predefined threshold $p$. Rather than always selecting the most likely image, this method favors high-probability outcomes while allowing for variation in the generated sequence.
@@
To find the optimal hyperparameters, the annotators compared pairs of beam search configurations to select the optimal beam width ($w$ in Eq. 5) and steps back ($m$ in Eq. 2). The winning configuration is compared against other configurations until a new configuration is selected as better. The selection criteria focused on semantic alignment and visual consistency. Semantic alignment assesses how well the image sequence matches the corresponding text, while visual consistency evaluates the coherence of the visuals across the sequence (e.g., keeping backgrounds, objects, or ingredients consistent).
Having chosen the optimal configuration, we compared BeamDiffusion to other methods. In the second annotation experiment, BeamDiffusion was compared with other baseline models. Annotators independently assessed the sequences, selecting the best method according to semantic and visual consistency with the possibility of choosing multiple options or none. For further information, see Appendix B.1.
Gemini Evaluation.
To further validate our findings and annotate a wider range of tasks, we used Gemini for both evaluation of configurations and baselines. Its selection process closely mirrored that of human annotators. For details, please refer to Appendix B.1.
Human-Gemini Agreement.
We conducted an alignment assessment to measure the agreement between human annotators and Gemini. For this, human annotators were asked to rate 16 Recipes, 10 DIYs, and 10 VIST tasks. These ratings were then compared to the evaluations made by Gemini. To quantify the agreement between the two, we computed Fleiss’ kappa [6], obtaining a value of 70%, which indicates a substantial level of agreement. Based on this strong alignment, we extended the evaluation to the remaining sequences using Gemini.
@@
Following prior work [2, 33, 21], we evaluate each method using: (1) Visual Consistency, based on average image similarity via DINO [27] (DINO-I) and CLIP [31] (CLIP-I); and (2) Semantic Alignment, using CLIP (CLIP-T) to measure similarity between the current image and the average of all step texts up to and including the current one. To place the scores on a comparable scale, we normalize CLIP-I, CLIP-T and DINO by their maximum. Finally, we compute CLIP* as the product between the CLIP-I and CLIP-T, and DINO* as the product between the DINO-I and CLIP-T, this formulations balances visual and semantic quality, while emphasizing semantic alignment, which is critical for accurately reflecting the stepwise input text. See Appendix B.1 for details.

### v3

# Latent Beam Diffusion Models for Generating Visual Sequences

@@ Abstract
While diffusion models excel at generating high-quality images from text prompts, they struggle with visual consistency when generating image sequences. Existing methods generate each image independently, leading to disjointed narratives — a challenge further exacerbated in non-linear storytelling, where scenes must connect beyond adjacent images. We introduce a novel beam search strategy for latent space exploration, enabling conditional generation of full image sequences with beam search decoding. In contrast to earlier methods that rely on fixed latent priors, our method dynamically samples past latents to search for an optimal sequence of latent representations, ensuring coherent visual transitions. As the latent denoising space is explored, the beam search graph is pruned with a cross-attention mechanism that efficiently scores search paths, prioritizing alignment with both textual prompts and visual context. Human and automatic evaluations confirm that BeamDiffusion outperforms other baseline methods, producing full sequences with superior coherence, visual continuity, and textual alignment.  [^1] Data and Code: huggingface.co/Gui28F/BeamDiffusion
@@ 1 Introduction
Image diffusion models [47, 11, 48] have made significant advancements in generating high-quality images. These models, especially when combined with text prompts, have shown great potential in producing detailed and contextually accurate visuals [39, 33, 42]. Despite the impressive results at generating individual images based on specific text prompts, diffusion models face challenges when it comes to creating coherent sequences of images. In most cases, each image is generated independently based on its own prompt, which does not naturally ensure visual continuity across a sequence of correlated prompts. This problem is further compounded when the narrative is non-linear, where a scene may be connected not only to the immediate prior scene but also to earlier scenes in the sequence.
Figure 1: BeamDiffusion leverages search in the latent denoising space to generate a coherent sequence of images from stepwise text instructions. The model conditions the generation of each new image $I_{j}$ on latents from previous steps $s_{j}$, forming a chain that carries forward information from past generations. At each step, multiple candidate latents are explored, the worst candidates are discarded, and the best ones are kept for the next step.
@@
We propose to research beam search as a scientifically well-grounded solution to this problem. Beam search is a core text generation technique for tasks like machine translation and standard inference [16, 54, 19, 41], allowing the model to explore multiple possible paths during the generation process, evaluating different trajectories and refining its predictions. However, despite its success in other domains, it has been barely explored in this context, leaving its potential untapped. In full image sequence generation, beam search could help navigate the complex latent space, refining coherence by adjusting to multiple text prompts or adding semantic cues, as illustrated in Figure 1.
@@
In this work, we revisit the latent diffusion framework with a focus on structured sequence generation. We aim to dynamizally sample the latent space of past diffusion processes and propose the BeamDiffusion method for improving multi-image visual consistency by leveraging shared latent representations in a principled way. BeamDiffusion explores the latent denoising trajectories with a beam search strategy for ensuring a conditional dependency between diffusion processes. As illustrated in Figure 2(a), our approach leverages a sampling strategy in the latent space to identify the best sequence of images, minimizing the risk of falling into suboptimal or incoherent paths, while maintaining visual continuity across images. As the beam diffusion algorithm traverses the latent space, the number of hypothesis grows and the search paths with the lowest probability are pruned to keep a fixed number of candidate paths. To this end, we incorporate a cross-attention mechanism [21] that evaluates and scores candidate image paths based on their alignment, providing a reliable quality measure for each path, allowing us to retain the best ones while pruning the least likely.
@@
To test the generalization and robustness of BeamDiffusion, we evaluated it in two diffusion architectures (LDM [39] and DiT [2]) and two domains (recipes and visual storytelling domains). Results obtained with human annotations, automatic evaluations and LLM-as-a-judge, all showed that the proposed approach outperforms baseline methods for generated sequence quality, showcasing its adaptability to diverse contexts.
@@ 3 Generating Non-Linear Visual Sequences
Generating a full sequence of visually coherent images from textual scene descriptions is a challenging research problem. Given a sequence of steps $S=\{s_{1},s_{2},\dots,s_{L}\}$, where each sequence step $s_{j}$ is a textual description, our goal is to generate a corresponding sequence of images $\{I_{1},I_{2},\dots,I_{L}\}$ that maintain consistency across all steps of the sequence. Diffusion Models generate high-quality images in a compressed latent space, but typically treat each image independently from a randomly initialized latent $z_{T}$.
Within this framework, Latent Diffusion Models (LDMs) [39] and Diffusion Transformers (DiTs) [32] share a common structure: both encode an input image $I$ into a latent representation $z$ using an encoder $\mathcal{E}$ and reconstruct it through a decoder $\mathcal{D}$ [39]. The reverse diffusion process is performed by a denoising backbone $\epsilon_{\theta}$, which can be a U-Net [40] (for LDMs) or a transformer [32] (for DiT), and incorporates cross-attention to condition the denoising iterations on an external input, e.g., the output embeddings of a text encoder $\tau_{\theta}(s)$. Both models are trained to iteratively estimate a denoised latent with the loss:
$L_{\text{diff}}=\mathbb{E}_{\mathcal{E}(I),s,\epsilon\sim\mathcal{N}(0,1),t}\left[\|\epsilon-\epsilon_{\theta}(z_{t},t,\tau_{\theta}(s))\|_{2}^{2}\right]$ (1)
This formulation highlights that each denoising process is agnostic to other generations, making it challenging to maintain coherence across a sequence of related prompts. This is particularly important in non-linear narratives [46], where preserving character identity, lighting, and environmental details is crucial. To address these challenges, prior work [36, 23, 49, 3, 37] has proposed reusing latent denoising iterations from earlier steps, raising the key question: how can we determine the optimal set of latent states that maximizes coherence across different generations?
In the following subsections, we introduce BeamDiffusion (Sec. 3.1), detail latent space exploration with diffusion beams (Sec. 3.2), and outline our beam pruning strategy (Sec. 3.3).
@@ 3.1 Beam Diffusion Model
To address the limitations of standard diffusion models in generating coherent image sequences, we introduce the BeamDiffusion model. The proposed approach aims to enhance both the consistency and quality of image sequences by adopting a strategy inspired by beam search, but with a twist, it operates on latent denoising space. BeamDiffusion incorporates latent representations from previously generated images, Figure 2(a). These latent representations guide the generation of new candidates, ensuring that each new image aligns with the context of the evolving sequence. This approach allows visual elements from previous images to be carried over or referenced, maintaining contextual consistency and continuity, while also adapting to the next scene description.
In this article, the term step will refer to an element in the input sequence; and the term iteration will refer to one denoising iteration of the diffusion model.
3.2 Latent Space Exploration with Diffusion Beams
@@
Algorithm 1 BeamDiffusion Algorithm
1:  Input parameters: $\{s_{1},\ldots,s_{L}\}$: sequence of steps; $r$: number of random seeds; $m$: steps back; $w$: beam width.
2:  $\hat{B}_{1}\leftarrow\{\text{SD}(s_{1},rnd)\}^{r}$ $\triangleright$ Beam initialization with $r$ different images.
3:  for $j=2$ to $L$ do
4:   for all beam $b\in\hat{B}_{j-1}$ do
5:    $\mathcal{Z}_{j}\leftarrow$ $\{\hat{B}_{j-n},\ldots,\hat{B}_{j-1}\}$ $\triangleright$ Read denoising latents from previous steps.
6:    for all latent $z\in\mathcal{Z}_{j}$ do
7:     $I_{j}\leftarrow\text{SD}(s_{j},z)$ $\triangleright$ Generate candidate a image.
8:     $\hat{B}_{j-1}\leftarrow\hat{B}_{j-1}\bigcup{I_{j}}$ $\triangleright$ Extend beam with candidate image.
9:    end for
10:   end for
11:   $\text{score}(I_{j})\leftarrow\varphi(I_{j}\in\hat{B}_{j})$ $\triangleright$ Score the new image of the beam.
12:   $B_{j}\leftarrow\arg\max_{|B_{j}|=w}\text{score}(\hat{B}_{j})$ $\triangleright$ Prune beams and keep top $w$ beams.
13:  end for
14:  return best beam $B^{*}\leftarrow\arg\max_{B_{L}}\text{score}(B_{L})$
@@
Inspired by beam search [53], which maintains and extends the most promising candidates at each sequence step, we introduce diffusion beams to structure sequential image generation. At each sequence step $j$, the set of candidate images $\mathcal{I}_{j}$ is used to compute the set of candidate diffusion beams $\hat{B}_{j}$: for each step $j$ we consider all the beams existing in step j1j-1 and extend each one of such $b$ beams $\hat{B}_{j-1}^{(b)}$ by pairing it with a new candidate image $I_{j}^{(k)}\in\mathcal{I}_{j}$. The resulting beams are defined as
@@
3.3 Pruning Diffusion Beam Paths
@@
The contrastive classifier $\varphi(I_{j}|\{s_{1},\ldots,s_{j}\},\{I_{1},\ldots,I_{j-1}\})$ is trained as a next image prediction task to estimate how well aligned an image is with the previously sequence of steps and images. For a given sequence $t$ and sequence step $j$, the ground-truth label $l_{t,j}$ indicates whether image $I_{j}$ belongs to the sequence given its prompt $s_{j}$ and context $B_{j-1}^{(b)}$. Negative examples are hallucinated from other sequences. The model minimizes the cross-entropy loss encouraging the model to promote candidates that best match the sequence’s semantic and visual dynamics. The classifier is trained on the Recipes [3] and VIST [15] datasets. See appendix H for implementation details.
@@ 4 Experimental Setup
In this section, we describe the experimental setup for evaluating BeamDiffusion’s performance in generating coherent image sequences. We first outline the different decoding strategies that are considered. We then present the datasets and baseline methods used, and describe the evaluation protocols, which include human judgments, automatic metrics, and LLM-as-a-judge evaluations using Gemini 2.5 [6], all designed to assess semantic and visual consistency across generated sequences.
Decoding strategies.
Our proposed method, BeamDiffusion, uses a beam search over latent sequences to generate coherent image sequences from textual prompts. To evaluate the effectiveness of this decoding strategy, we also compare it against two alternative approaches applied to the same model: Greedy / CoSeDlen=1 [37], which selects at each step the top-scoring image according to CLIP similarity from the first few latents of the previous step, focusing on immediate best-match selection; and Nucleus Sampling [13], a probabilistic strategy that samples images from the subset whose cumulative CLIP probability exceeds a threshold $p$, encouraging diversity while still prioritizing high-probability outcomes. All strategies are applied to the same sequences of prompts using two underlying diffusion backbones, FLUX.1-dev [2] and SD2.1 [39], allowing a direct comparison of their ability to maintain sequence-level coherence across different models.
@@
We used publicly available manual tasks in the Recipes domain [3], along with a small portion of the Visual Storytelling (VIST) dataset [15], using an equal number of tasks from each. Recipes provide detailed, step-by-step instructions, while the sampled VIST tasks offer additional diversity for training the classifier. This combination allows us to evaluate BeamDiffusion’s performance on both familiar and novel task types. Together, these datasets enable a thorough assessment of sequence coherence, semantic alignment, and visual consistency. More details can be found in Appendix F.
@@
To assess the performance of BeamDiffusion, we compared against several competitive baselines that aim to generate coherent image sequences from text: GILL [20], StackedDiffusion [29], StoryDiffusion [55], and One-Prompt-One-Story (1P1S) [24].
@@
To find the optimal hyperparameters, the annotators compared pairs of beam search configurations to select the optimal beam width ($w$ in Eq. 5) and steps back ($m$ in Eq. 2). The winning configuration is selected iteratively by comparing different configurations and preserving always the better one. See Appendix B for the detailed results.
Having chosen the optimal configuration, we compared BeamDiffusion to other methods. In the second annotation experiment, BeamDiffusion was compared with other baseline models. Annotators independently assessed the sequences, selecting the best method according to semantic and visual consistency with the possibility of choosing multiple options or none. For further information, see Appendix E.
LLM-as-a-judge Evaluation.
To further validate our findings and annotate a wider range of tasks, we adopt the LLM-as-a-judge paradigm, leveraging Gemini, to evaluate both configurations and baselines. Its selection process closely mirrored that of human annotators. For details, please refer to Appendix E.
Human and LLM-as-a-judge Agreement.
We conducted an alignment assessment to measure agreement between human annotators and the LLM-as-a-judge. Seven annotators rated 16 Recipes and 10 VIST tasks, which we compared to the model’s evaluations. To quantify the agreement between the two, we computed Fleiss’ kappa [8], obtaining a value of 70%, which indicates a substantial level of agreement. Building on this strong alignment, we employed the LLM-as-a-judge to evaluate the remaining sequences.
@@
Following prior work [3, 37, 23], we evaluate each method using two complementary criteria: (1) Visual Consistency, measured via average image similarity using DINO [30] (DINO-I) and CLIP [35] (CLIP-I); and (2) Semantic Alignment, computed with CLIP (CLIP-T) by comparing each image to the average of all step texts up to and including that step. To ensure comparability across metrics, we normalize CLIP-I, CLIP-T, and DINO-I by their respective maxima. We further define CLIP* as the product of CLIP-I and CLIP-T, and DINO* as the product of DINO-I and CLIP-T, balancing visual and semantic quality while emphasizing alignment with the stepwise input text. To better capture real performance, the results for DINO-I, CLIP-I, CLIP-T, CLIP*, and DINO* are reported as win percentages rather than mean values. This avoids overestimating methods that perform very well on some sequences but poorly on others, which could otherwise produce artificially high averages despite inconsistent quality.
In addition, we adopt metrics from StackedDiffusion [29] to evaluate task-specific aspects of the generated sequences: Goal Faithfulness, measuring alignment of the final image with the overall textual goal; Step Faithfulness, measuring alignment of each intermediate image with its corresponding step text; and Cross-Image Consistency, capturing visual coherence across images in the sequence. These metrics complement the previous ones by explicitly assessing semantic fidelity and sequence-level consistency. See Appendix E for further details.